Grade 12Engineering Entrance Exams

The bulk modulus of a gas is 6 × 103 N/m2. the additional pressure needed to reduce the volume of the gas by 10% is (a) 300 N/m2 (b) 400 N/m2 (c) 1000 N/m2 (d) 600 N/m2

Abhishek
12 Years agoGrade 12

2 Answers

Saurabh Koranglekar
6 Years ago
Dear student

The answer is option d

Regards
Vikas TU
6 Years ago
Bulk modulus B = 6 × 10³ N / m²
= > Δ V / V × 100 = 10 %
= > Δ V / V = 10 / 100
= > Δ V / V = 1 / 10
= > Δ V / V = 10⁻¹
We know :
B = P / ( Δ V / V )

= > P = B × Δ V / V
Putting values here we get :
= > P = 6 × 10³ × 10⁻¹
= > P = 6 × 10²
= > P = 600 N / m²
 

question mark
Hello,
 
Can we learn and prepare for MHT-CET online ?
engineering entrance exams

Last Activity: 4 Years ago